• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name J. Wexler

    Log entry time 08:00:38 on August27,2009

    Entry number 286225

    keyword=Shift Summary - Owl 8/27

    Shift summary for the owl shift, 8/27

    12AM - No beam, still in controlled access.
    2:20 - DAQ issue solved on the Moller expt., going to beam permit.
    2:50 - An issue arose with the one of the magnets. We request access to assess the issue.
    3:20 - Magnets still broken, expert is called in to fix magnet.
    4:20 - The magnet is repaired, we're going to beam permit and ramping the spectrometer magnets.
    5:30 - We're preparing the thick tantalum target, going through the MCC plan to give us 60+ uA on the target in order to test lumi noise, and then check the corners of a 2mm x 2mm area.
    6:00 - Multiple changes between empty, tantalum, and BeO targets occured in the MCC plan to get the current.
    6:20 - Started the putting 10 uA beam on thick tantalum.
    6:35 - Up to 70 uA. Begin looking at lumi noise
    6:50 - In the name of letting the target cool, we've gone back to 5 uA.
    7:15 - We called the MCC, asking if they can ramp the current above 70 uA in order to see the effect on lumi noise.
    7:25 - MCC can't stably reach above 65 uA due to issues on their end. We reduce the current to 10uA.